OpenStudy (anonymous):

write the expression log (sub 2) 7 - log (sub 2) 9 + 2log (sub 2) x as a single algorithm

OpenStudy (anonymous):

\[log_2(7)-log_2(9)+2log_2(x)\] So we need to know three Logarithm rules: \(log_a(x)+log_a(y)=log_a(xy)\) \(log_a(x)-log_a(y)=log_a\left(\frac{x}{y}\right)\) \(y \phantom{.} log_a(x)=log_a(x^y)\) So we can proceed: \(log_2(7)-log_2(9)+2log_2(x)\) \(=log_2\left(\frac{7}{9}\right)+log_2(x^2)\) \(=log_2\left(\frac{7x^2}{9}\right)\)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np